# Observation of the Galactic Center PeVatron Beyond 100 TeV with HAWC

### v1

@@ 3 Main analysis results
where is the flux normalization at the pivot energy and is the power-law index. The pivot energy of 26 TeV is calculated such that it minimizes the correlation between the flux normalization and spectral index. We summarize the best-fit parameters of HAWC J1746-2856 in Table 1 and include both statistical and systematic uncertainties. The latter account for the contribution of four non-negligible independent systematic uncertainties that were identified in the previous energy-dependent study of the Crab (Abeysekara et al., 2019) and are estimated by producing instrument response functions with different detector configurations to investigate any potential mis-modeling of the detector. The results were then compared with the standard HAWC analysis and the uncertainties were added in quadrature. We do not model the diffuse background emission because the flux detected by HAWC from the GC is much higher than the expected Galactic diffuse emission (Cao et al., 2023; Alfaro et al., 2024).
We calculated an UL on the minimum energy at 6 TeV and a lower limit on the maximum energy at 114 TeV, both at the 68% CL. Above 100 TeV, 3474 events passed trigger conditions for reconstruction, from which 98 events passed HAWC gamma/hadron separation cuts. At 100 TeV, the energy resolution is 10% in (Abeysekara et al., 2019) and the hadron retention after gamma/hadron separation cuts is %. To count the events, we used a circular region centered in the best-fit position (see Table 1) with the radius set at the UL on the source extension.
@@
In Figure 1(d), we subtract the flux from the two H.E.S.S. point sources from the HAWC best-fit spectrum (shown separately in Figure 1(b)). The error band illustrates the combination of HAWC and H.E.S.S. uncertainties in quadrature. We also compare our measurement with the diffuse emission flux points estimated in H.E.S.S. Collaboration et al. (2016), where the diffuse emission was derived within an annulus of inner radius 0.15∘—to exclude HESS J1745-290—and outer radius 0.45∘. In that study, a sector (66∘) of the annulus is excluded to avoid HESS J1746-285. These excluded regions may explain the slightly larger flux detected by HAWC, although the uncertainties from both results are still compatible (see Figure 1(d)). Other IACTs have also measured the diffuse emission, however, they use regions with significantly different morphologies (Acciari et al., 2020; Adams et al., 2021).

### v2

@@ 3 Main analysis results
where is the flux normalization at the pivot energy and is the power-law index. The pivot energy of 26 TeV is calculated such that it minimizes the correlation between the flux normalization and spectral index. We summarize the best-fit parameters of HAWC J1746-2856 in Table 1 and include both statistical and systematic uncertainties. The latter account for the contribution of four non-negligible independent systematic uncertainties that were identified in the previous energy-dependent study of the Crab (Abeysekara et al., 2019) and are estimated by producing instrument response functions (IRFs) with different detector configurations to investigate any potential mis-modeling of the detector. The results were then compared with the standard HAWC analysis and the uncertainties were added in quadrature. Another source of systematic uncertainty in the flux of HAWC J1746-2856 could be emission from background cosmic rays, often referred to as the cosmic-ray sea, which is thought to have a consistent energy density throughout the Galaxy. Locally, above 100 TeV, the energy density of the cosmic-ray spectrum is approximately eV/cm3 (Aguilar et al., 2015). For the diffuse emission to significantly impact the results, one would need to assume that the cosmic-ray sea’s flux is nearly a factor of ten higher at the Galactic Center. Even assuming an unusual spectral dependence or normalization of the diffusion coefficient within the CMZ, it would be difficult to explain the at least tenfold discrepancy between the reported local spectrum and the CMZ spectrum observed by HAWC. Additionally, the spectral indices of the cosmic-ray sea and HAWC J1746-2856 are not compatible. Thus, we keep the simplest point-source model.
We calculated an UL on the minimum energy at 6 TeV and a lower limit on the maximum energy at 114 TeV, both at the 68% CL. Above 100 TeV, the significance of the signal is 1.2 . Above 100 TeV, 3474 events passed trigger conditions for reconstruction, from which 98 events passed HAWC gamma/hadron separation cuts. At 100 TeV, the energy resolution is 10% in (Abeysekara et al., 2019) and the hadron retention after gamma/hadron separation cuts is %. To count the events, we used a circular region centered in the best-fit position (see Table 1) with the radius set at the UL on the source extension.
@@
In Figure 1(d), we subtract the flux from the two H.E.S.S. point sources from the HAWC best-fit spectrum (shown separately in Figure 1(b)). The error band illustrates the combination of HAWC and H.E.S.S. uncertainties in quadrature. We also compare our measurement with the diffuse emission flux points estimated in H.E.S.S. Collaboration et al. (2016), where the diffuse emission was derived within an annulus of inner radius 0.15∘—to exclude HESS J1745-290—and outer radius 0.45∘. In that study, a sector (66∘) of the annulus is excluded to avoid HESS J1746-285. These excluded regions and the slightly larger radius of the HAWC source (0.46∘) may explain the higher flux detected by HAWC, although both results are still compatible within uncertainties (see Figure 1(d)). The hard spectrum reported by H.E.S.S. with a photon index of 2.3 (H.E.S.S. Collaboration et al., 2016) is mostly dominated by events with energies below 10 TeV. HAWC is more sensitive at higher energies and measures an index of 2.9. The change in the spectral index occurs at low energies where HAWC is not sensitive enough to probe the cause, given the large zenith angle. However, we find no evidence of significant spectral curvature from 10s of TeV to 114 TeV. Other IACTs have also measured the diffuse emission. However, they use regions with significantly different morphologies: in the studies by H.E.S.S. (Abdalla et al., 2018) and MAGIC (Acciari et al., 2020) the entire l GC region is included, while VERITAS Adams et al. (2021) utilized seven circular regions of 0.1° radius outside of the H.E.S.S. (H.E.S.S. Collaboration et al., 2016) annuli.